OpenStudy (anonymous):

Using the fact that lim h->0 sin(h)/h = 1 and lim h->0 cos(h)-1/h =0 compute the following limits a. lim h->0 sin(x+h)-sin(x)/h b lim h->0 cos(x+h)-cos(x)/h

OpenStudy (anonymous):

addition angle formula for the first one second one too

OpenStudy (anonymous):

\[\sin(x+h)-\sin(x)=\sin(x)\cos(h)+\sin(h)\cos(x)-\sin(x)\] divide all by \(h\) and take the limit

OpenStudy (anonymous):

you get that?

OpenStudy (anonymous):

I think how about b

OpenStudy (anonymous):

exactly the same, but this time the addition angle formula for cosine

OpenStudy (anonymous):

i can walk you through one if you like the other is the same

OpenStudy (solomonzelman):

satelline, you will do all of that work?

OpenStudy (anonymous):

btw do you know what the answer is before you begin? @SolomonZelman what work?

OpenStudy (anonymous):

k please

OpenStudy (solomonzelman):

to which one ? Also, I am logging off in 2 minutes... SCHEDULE :P

OpenStudy (solomonzelman):

sorry

OpenStudy (anonymous):

lets do the first one

OpenStudy (anonymous):

k

OpenStudy (solomonzelman):

I am not a good teacher anyway ... bye

OpenStudy (anonymous):

better than me

OpenStudy (anonymous):

before we begin, is it clear that you are being asked for the derivative of sine, and that the answer is cosine? to the first one i mean

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

k lets ignore the limit for the moment so i don't have to write it in, and just to the algebra first

OpenStudy (anonymous):

\[\frac{\sin(x+h)-\sin(x)}{h}=\frac{\sin(x)\cos(h)+\sin(h)\cos(x)-\sin(x)}{h}\] is a start that by the "addition angle formula for sine

OpenStudy (anonymous):

now break in to two parts \[\frac{\sin(x)\cos(h)-\sin(x)}{h}+\frac{\sin(h)\cos(x)}{h}\] factor the first one and get \[\sin(x)\frac{\cos(h)-1)}{h}+\frac{\sin(h)\cos(x)}{h}\]

OpenStudy (anonymous):

now take the limit as \(h\to 0\)

OpenStudy (anonymous):

since \[\lim_{h\to 0}\frac{\cos(h)-1}{h}=0\] the first term goes bye bye and since \[\lim_{h\to 0}\frac{\sin(h)}{h}=1\] the second term becomes \(\cos(x)\) and needed

OpenStudy (anonymous):

I think I got it but what formula for b?

OpenStudy (anonymous):

the addition angle for cosine is \[\cos(x+h)=\cos(x)\cos(h)-\sin(x)\sin(h)\]

OpenStudy (anonymous):

do you have time to show the work?

OpenStudy (anonymous):

i have the time, but why don't you try it first if you mimic what i did above, you will a) understand it better and b) learn what it is about it you don't understand then ask

OpenStudy (anonymous):

k thanks so much

OpenStudy (anonymous):

yw write if you get stuck